# ExoJAX Retrievals of VLT/CRIRES Spectra of Luhman 16AB: C/O Ratios and Systematic Uncertainties

### v1

@@ 4.2.1 Comparison with Medium-Resolution Flux Density
4.2.2 Impact of Photometric Variability in Luhman 16B
@@
Luhman 16A shows $2\%$ variability over a 7-hr baseline (Biller2024), and Luhman 16B exhibits $\sim$3% variability over 5 hr (Crossfield2014) and up to $\sim$6% over 7 hr (Biller2024). Such variability, common among L/T transition brown dwarfs including Luhman 16AB, is often attributed to heterogeneous silicate/iron clouds whose evolving rotational modulation produces temporal brightness changes (Radigan2014; Crossfield2014; Chen2024; Saumon_Marley2008). The CRIRES time series used here comprise 14 spectra ($\sim$5 hr) for each component (Crossfield2014) and show comparable normalized flux variability: 2.3% for A and 3.2% for B.
@@
Our baseline retrieval assumes a horizontally homogeneous (i.e., effectively averaged) atmosphere, neglecting cloud patchiness. To assess potential systematics on C/O and cloud‐top pressure from this assumption, we performed independent retrievals for each epoch of Luhman 16B using $\mathrm{HITEMP_{H_{2}}\,\,CO}$. We find no statistically significant differences relative to the stacked‐spectrum results in C/O or cloud deck pressure (Figure 7); quantitatively, the epoch-by-epoch C/O values agree with the stacked value ($\mathrm{HITEMP_{H_{2}}\,\,CO}$) to within at most $1.3\sigma$, and the cloud deck pressures to within at most $2.6\sigma$. Hence, for the present data, the impact of ignoring cloud inhomogeneity appears subdominant to systematics associated with the line lists. This conclusion is consistent with the analysis of the $J$-band VLT/CRIRES+ spectrum by deRegt2025, who found that a patchy-cloud model is only weakly preferred for Luhman 16B and that the retrieval outcomes are largely insensitive to the inclusion of patchiness.
@@
Figure 7: C/O and cloud deck pressure for Luhman 16B from stacked and time-resolved spectra. Left: C/O. Right: cloud deck pressure. In each panel, the region left of the vertical dashed line shows results from stacked spectra using three CO line lists (orange: $\mathrm{ExoMol}$; green: $\mathrm{HITEMP_{Air}}$; blue: $\mathrm{HITEMP_{H_{2}}}$). The region right of the dashed line shows epoch-by-epoch (time-resolved) retrievals using $\mathrm{HITEMP_{H_{2}}\,\,CO}$; from the 14 spectra spanning 0.0–5.0 hr we use five representative epochs. Error bars denote 90% credible region. In the left panel, the dotted horizontal line marks the solar C/O (=0.59=0.59; Asplund2021).
5 Retrievals Using a Flexible Gaussian Process TTPP Profile
@@
Some high-resolution retrieval studies of brown-dwarf and exoplanet atmospheres adopt simple parameterizations of the TTPP profile, such as a power-law form (e.g., Kawahara2022; Kawashima2025; Kasagi2025). Although efficient, such profiles can be overly restrictive. More flexible parameterizations have therefore been proposed, including schemes that parameterize the temperature gradient in discrete pressure layers (Zhang2023) and $P$-splines (e.g., Line2015; deRegt2025). However, these approaches still remain tied to specific functional forms and may fail to capture complex structures.
@@
As a potential future extension that takes advantage of the computational efficiency of HMC/NUTS in high-dimensional settings, [^1] The convergence time of conventional random Metropolis–Hastings scales as $D^{2}$ with $D$ being the number of parameters, whereas HMC scales as $D^{5/4}$ (HMC). There are also studies suggesting that NUTS scales as $D^{1/4}$ when initialized in the concentration region of the canonical Gaussian measure (Bou-Rabee2024). Kawahara2022 proposed a nonparametric TTPP profile model based on Gaussian processes (GP, 2023ARA&A..61..329A). That work, however, did not explore its practical applicability in real retrievals, its sensitivity to other parameters, or its validation through injection–recovery tests. Here we refine the GP-based TTPP profile model into a more practical framework and demonstrate its performance through application to the retrieval of Luhman 16AB data.
5.1 Formulation of the GP TTPP Profile
@@
Relative to the GP formulation of Kawahara2022, our implementation introduces several changes aimed at practical convergence and physical regularization. First, while Kawahara2022 adopted a RBF kernel, we instead use a Matérn 3/2 kernel. The Matérn kernel is less smooth than the RBF kernel, avoiding the overly smooth behavior inherent to the RBF assumption. We also infer the kernel hyperparameters (the amplitude $a_{u}$ and correlation length $\tau_{u}$) jointly with the atmospheric parameters, thereby removing the need to set them manually. In addition, instead of adopting a constant mean function as in Kawahara2022, we use a power-law form characterized by $(T_{0,\,m},\,\alpha_{m})$, whose parameters are inferred jointly with the GP. We find that a constant mean tends to force the retrieved profile to flatten toward the average value in the upper and lower layers, as is also hinted in figure 11 Kawahara2022. The power-law mean model, on the other hand, better reflects the physically expected structure of self-luminous brown dwarfs while allowing the GP to capture deviations from it.
5.2 Demonstration
@@ 5.3 Application to Luhman 16AB
Figure 13: Comparison between the retrieved volume mixing ratios and FastChem equilibrium predictions (solar abundances from Asplund2021). (Left) Luhman 16A. (Right) Luhman 16B.
We applied the GP TTPP retrieval to the CRIRES data of Luhman 16AB. Best-fit model spectra for both components are shown in Figure 10. Corner plots for the selected parameters are shown in Figure 11; the retrieved TTPP profiles are shown in Figure 12 along with the power-law profiles from Section 4; the retrieved parameter values are summarized in the rightmost columns of Tables 2 and 3.
@@ 6 Summary and Discussion
6.1 Comparison with Previous Work
@@
The molecules included in our model, H2O, CH4, CO, and HF, have all been detected in the $H$- and $K$-band Gemini South/IGRINS high-resolution ($R\sim 28,000$) spectra of Luhman 16AB presented by Ishikawa2025. Their study compared the spectra against forward models generated with PICASO, which incorporates radiative–convective equilibrium, Sonora Bobcat TTPP profiles, VIRGA cloud models, and nonequilibrium chemistry parameterized by a vertical eddy diffusion coefficient $K_{zz}$. The retrieved volume mixing ratios (VMRs) of the common molecules (their Figure 8) are broadly consistent with ours, but tend to be smaller, likely because they fixed the metallicity to solar whereas we inferred slightly super-solar values. Both H2O and HF have also been detected in the CRIRES+/VLT $J$-band high-resolution retrieval by deRegt2025. Their modeling adopted a TTPP parameterization based on the temperature gradient at five knots (as in Zhang2023), and clouds were treated as gray opacity decreasing with a power-law of pressure from the cloud base upward. Despite these differences, the retrieved molecular abundances are consistent with ours. Neither study provides explicit estimates of the C/O ratio: Ishikawa2025 fixed the metallicity and C/O to solar, and in the spectral range analyzed by deRegt2025, no carbon-bearing molecules were detected.
@@
Ishikawa2025 also detected H2 lines. By comparing the inferred H2 column density with the interstellar gas-to-dust ratio, they suggested that dust clouds may reside deeper than the H2 absorption layer, located at $\sim 3\,\mathrm{bar}$ (or $\log_{10}P\sim 0.5$) in both components A and B. Their self-consistent models likewise imply that clouds are present at similar or deeper pressures relative to the H2 line-forming region. In the power-law cloud model of deRegt2025, the cloud opacity is dominated by contributions around $\log_{10}P\sim 0.5$–1, and this estimate is robust to assumptions about patchiness. In our analysis as well, the inferred cloud top pressures consistently fall within $\log_{10}P\sim 0.5$–1, broadly agreeing with these results. On the other hand, the quantitative estimates depend on assumptions about the TTPP profile and line lists, which in turn propagate into the inferred C/O ratio as discussed in the next subsection.
6.2 $\mathrm{C/O}$ Ratios
@@
Figure 14 compares $\mathrm{C/O}$ inferred with each CO line list configuration and with the GP TTPP model (using $\mathrm{HITEMP_{H_{2}}\,\,CO}$) to literature values for other brown dwarfs and super Jupiters. For both A and B, the choice of CO line list introduces a systematic offset of $\sim$7% in $\mathrm{C/O}$. By contrast, the assumed TTPP profile (power-law versus GP) does not yield a significant difference in $\mathrm{C/O}$. An epoch-by-epoch analysis of the Luhman 16B time series likewise reveals no statistically significant impact from cloud-cover variability on the retrieved $\mathrm{C/O}$ or molecular VMRs (Section 4.2.2). We therefore conclude that, in our framework, the principal systematic on $\mathrm{C/O}$ arises from the CO line list, whereas the TTPP shape and photometric variability are not dominant. In all cases, the C/O values for components A and B agree within $\approx 2\sigma$, consistent with the expectation that isolated brown dwarfs (and their binaries) form in a star-like manner.
@@
At face value, Figure 14 hints that brown dwarfs may exhibit higher $\mathrm{C/O}$ than super Jupiters. Although this is a collection of inhomogeneous inferences from the literature, our measurements for Luhman 16AB, together with the associated systematic-error budget, reinforce this possibility. While differences in formation pathways (i.e., gravitational instability vs. core accretion) remain a possible explanation, the contrast could also reflect differences in the efficiency of oxygen sequestration by silicate (and other) clouds, as emphasized in previous low-resolution studies (e.g., Line2015; Line2017): more efficient rainout in field brown dwarfs would raise the gas-phase C/O, whereas weaker sequestration in super Jupiters would leave C/O closer to the stellar value. A decisive test will require a larger, uniformly analyzed sample of both planets and brown dwarfs, so that inter-study systematics are minimized and C/O can be compared on equal footing. Building on such a uniform baseline, joint analyses of high-resolution spectra (to constrain TTPP) and JWST coverage (to probe silicate or other cloud composition and vertical structure) can further mitigate cloud–TTPP–chemistry degeneracies, thereby providing a more direct test of the efficiency of oxygen sequestration by rainout. We return to the prospects and implementation of such joint retrievals in Section 6.4.
@@
6.3 Usefulness and Caveats of the GP TTPP Model
For both simulations and real data, the GP TTPP model captures structures that are more complex than a simple power-law and yields a more faithful assessment of uncertainties. In the present application, these differences have only a modest impact on bulk parameters such as $\mathrm{C/O}$. However, as discussed in Section 5.3, comparisons between retrieved molecular abundances and chemical-equilibrium predictions are sensitive to how accurately TTPP uncertainties are estimated; in this context, the GP framework is important because it enables predictive uncertainties to faithfully reflect the spectra’s information content, remaining conservative where the data are weakly informative. The GP approach is also sensitive to residual model–data mismatches and thus shows promise for diagnosing deficiencies in current models, potentially pointing to additional physics.
On the other hand, as indicated in Section 5.3, degeneracies between the TTPP shape and cloud opacity can also allow nonphysical temperature structures if left unconstrained. Mitigations include imposing weak physical regularization (e.g., monotonicity or stability constraints) on flexible TTPP profiles and/or extending the analysis to broader wavelength coverage where cloud signatures are stronger. The latter strategy can help balance flexibility with physical plausibility; see also Section 6.4.
Figure 15: Comparison between the medium-resolution spectra predicted by extrapolating our high-resolution retrievals and the JWST observations. Shaded bands show the 90% credible intervals of the predicted spectra computed with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a power-law TTPP profile (blue: Luhman 16A; orange: Luhman 16B). Points show the corresponding JWST medium-resolution spectra from Biller2024; Chen2025 (blue: A; orange: B).
6.4 Cloud Degeneracies and Joint Analysis with JWST
@@
We gratefully acknowledge Ian J. M. Crossfield for providing the reduced VLT/CRIRES spectra of Luhman 16AB. Based on observations collected at the European Southern Observatory under ESO programme(s) 291.C-5006(A). We also thank the anonymous reviewer for providing comments that helped improve the manuscript. This study was supported by JSPS KAKENHI Grant Numbers 24KJ1555 (H.Y.), 21H04998 (K.M., Y.K., and H.K.), 22H05150 (Y.K.), 25K17429 (Y.K.), 23H01224 (Y.K. and H.K.), and 23H00133 (H.K.). We used ChatGPT (OpenAI, GPT-5, Sept 2025) for language editing. No part of the scientific analysis was performed by ChatGPT. The authors are responsible for the content and interpretation. Table 4 provides the literature values, together with the corresponding references, for the objects shown in Figure 1.
@@
Gl 229 B FBD Y $0.77\pm 0.01$ $71.47^{+0.57}_{-0.61}$ccGl 229 B has been reported to be a binary, based on high–angular-resolution observations with GRAVITY/VLTI together with radial-velocity monitoring with CRIRES+/VLT (Xuan2024b). $850\pm 50$ (6), (7)
@@
Note. — References—(1) deRegt2024; (2) Xuan2022; (3) Xuan2024c; (4) Hsu2024; (5) Currie2020; (6) Kawashima2025; (7) Calamari2022; (8) Wang2022; (9) Mulder2025; (10) Schlawin2017; (11) Zapatero2014; (12) Filippazzo2015; (13) Costes2024; (14) Franson2022; (15) Johnson-Groh2017; (16) Picos2024; (17) Scholz2005; (18) Cooper2024; (19) Morales2021; (20) Zhang2021; (21) Vos2022; (22) Zhang2024b; (23) Currie2023; (24) Landman2024; (25) Lacour2021; (26) Chilcote2017; (27) Picos2025; (28) Schwarz2016; (29) Zhang2024a; (30) Wood2023; (31) Xuan2024a; (32) Gandhi2025; (33) Bonnefoy2010

### v2

@@ IV.2.1 Comparison with Medium-Resolution Flux Density
IV.2.2 Impact of Photometric Variability in Luhman 16B
@@
Luhman 16A shows $2\%$ variability over a 7-hr baseline (Biller et al., 2024), and Luhman 16B exhibits $\sim$3% variability over 5 hr (Crossfield et al., 2014) and up to $\sim$6% over 7 hr (Biller et al., 2024). Such variability, common among L/T transition brown dwarfs including Luhman 16AB, is often attributed to heterogeneous silicate/iron clouds whose evolving rotational modulation produces temporal brightness changes (Radigan, 2014; Crossfield et al., 2014; Chen et al., 2024; Saumon and Marley, 2008). The CRIRES time series used here comprise 14 spectra ($\sim$5 hr) for each component (Crossfield et al., 2014) and show comparable normalized flux variability: 2.3% for A and 3.2% for B.
@@
Our baseline retrieval assumes a horizontally homogeneous (i.e., effectively averaged) atmosphere, neglecting cloud patchiness. To assess potential systematics on C/O and cloud‐top pressure from this assumption, we performed independent retrievals for each epoch of Luhman 16B using $\mathrm{HITEMP_{H_{2}}\,\,CO}$. We find no statistically significant differences relative to the stacked‐spectrum results in C/O or cloud deck pressure (Figure 7); quantitatively, the epoch-by-epoch C/O values agree with the stacked value ($\mathrm{HITEMP_{H_{2}}\,\,CO}$) to within at most $1.3\sigma$, and the cloud deck pressures to within at most $2.6\sigma$. Hence, for the present data, the impact of ignoring cloud inhomogeneity appears subdominant to systematics associated with the line lists. This conclusion is consistent with the analysis of the $J$-band VLT/CRIRES+ spectrum by de Regt et al. (2025), who found that a patchy-cloud model is only weakly preferred for Luhman 16B and that the retrieval outcomes are largely insensitive to the inclusion of patchiness.
@@
Figure 7: C/O and cloud deck pressure for Luhman 16B from stacked and time-resolved spectra. Left: C/O. Right: cloud deck pressure. In each panel, the region left of the vertical dashed line shows results from stacked spectra using three CO line lists (orange: $\mathrm{ExoMol}$; green: $\mathrm{HITEMP_{Air}}$; blue: $\mathrm{HITEMP_{H_{2}}}$). The region right of the dashed line shows epoch-by-epoch (time-resolved) retrievals using $\mathrm{HITEMP_{H_{2}}\,\,CO}$; from the 14 spectra spanning 0.0–5.0 hr we use five representative epochs. Error bars denote 90% credible region. In the left panel, the dotted horizontal line marks the solar C/O (=0.59=0.59; Asplund et al., 2021).
V Retrievals Using a Flexible Gaussian Process TTPP Profile
@@
Some high-resolution retrieval studies of brown-dwarf and exoplanet atmospheres adopt simple parameterizations of the TTPP profile, such as a power-law form (e.g., Kawahara et al., 2022; Kawashima et al., 2025; Kasagi et al., 2025). Although efficient, such profiles can be overly restrictive. More flexible parameterizations have therefore been proposed, including schemes that parameterize the temperature gradient in discrete pressure layers (Zhang et al., 2023) and $P$-splines (e.g., Line et al., 2015; de Regt et al., 2025). However, these approaches still remain tied to specific functional forms and may fail to capture complex structures.
@@
As a potential future extension that takes advantage of the computational efficiency of HMC/NUTS in high-dimensional settings, [^1] The convergence time of conventional random Metropolis–Hastings scales as $D^{2}$ with $D$ being the number of parameters, whereas HMC scales as $D^{5/4}$ (Neal, 2011). There are also studies suggesting that NUTS scales as $D^{1/4}$ when initialized in the concentration region of the canonical Gaussian measure (Bou-Rabee and Oberdörster, 2024). Kawahara et al. (2022) proposed a nonparametric TTPP profile model based on Gaussian processes (GP, Aigrain and Foreman-Mackey, 2023). That work, however, did not explore its practical applicability in real retrievals, its sensitivity to other parameters, or its validation through injection–recovery tests. Here we refine the GP-based TTPP profile model into a more practical framework and demonstrate its performance through application to the retrieval of Luhman 16AB data.
V.1 Formulation of the GP TTPP Profile
@@
Relative to the GP formulation of Kawahara et al. (2022), our implementation introduces several changes aimed at practical convergence and physical regularization. First, while Kawahara et al. (2022) adopted a RBF kernel, we instead use a Matérn 3/2 kernel. The Matérn kernel is less smooth than the RBF kernel, avoiding the overly smooth behavior inherent to the RBF assumption. We also infer the kernel hyperparameters (the amplitude $a_{u}$ and correlation length $\tau_{u}$) jointly with the atmospheric parameters, thereby removing the need to set them manually. In addition, instead of adopting a constant mean function as in Kawahara et al. (2022), we use a power-law form characterized by $(T_{0,\,m},\,\alpha_{m})$, whose parameters are inferred jointly with the GP. We find that a constant mean tends to force the retrieved profile to flatten toward the average value in the upper and lower layers, as is also hinted in figure 11 Kawahara et al. (2022). The power-law mean model, on the other hand, better reflects the physically expected structure of self-luminous brown dwarfs while allowing the GP to capture deviations from it.
V.2 Demonstration
@@ V.3 Application to Luhman 16AB
Figure 13: Comparison between the retrieved volume mixing ratios and FastChem equilibrium predictions (solar abundances from Asplund et al. 2021). (Left) Luhman 16A. (Right) Luhman 16B.
We applied the GP TTPP retrieval to the CRIRES data of Luhman 16AB. Best-fit model spectra for both components are shown in Figure 10. Corner plots for the selected parameters are shown in Figure 11; the retrieved TTPP profiles are shown in Figure 12 along with the power-law profiles from Section IV; the retrieved parameter values are summarized in the rightmost columns of Tables 2 and 3.
@@ VI Summary and Discussion
VI.1 Comparison with Previous Work
@@
The molecules included in our model, H2O, CH4, CO, and HF, have all been detected in the $H$- and $K$-band Gemini South/IGRINS high-resolution ($R\sim 28,000$) spectra of Luhman 16AB presented by Ishikawa et al. (2025). Their study compared the spectra against forward models generated with PICASO, which incorporates radiative–convective equilibrium, Sonora Bobcat TTPP profiles, VIRGA cloud models, and nonequilibrium chemistry parameterized by a vertical eddy diffusion coefficient $K_{zz}$. The retrieved volume mixing ratios (VMRs) of the common molecules (their Figure 8) are broadly consistent with ours, but tend to be smaller, likely because they fixed the metallicity to solar whereas we inferred slightly super-solar values. Both H2O and HF have also been detected in the CRIRES+/VLT $J$-band high-resolution retrieval by de Regt et al. (2025). Their modeling adopted a TTPP parameterization based on the temperature gradient at five knots (as in Zhang et al., 2023), and clouds were treated as gray opacity decreasing with a power-law of pressure from the cloud base upward. Despite these differences, the retrieved molecular abundances are consistent with ours. Neither study provides explicit estimates of the C/O ratio: Ishikawa et al. (2025) fixed the metallicity and C/O to solar, and in the spectral range analyzed by de Regt et al. (2025), no carbon-bearing molecules were detected.
@@
Ishikawa et al. (2025) also detected H2 lines. By comparing the inferred H2 column density with the interstellar gas-to-dust ratio, they suggested that dust clouds may reside deeper than the H2 absorption layer, located at $\sim 3\,\mathrm{bar}$ (or $\log_{10}P\sim 0.5$) in both components A and B. Their self-consistent models likewise imply that clouds are present at similar or deeper pressures relative to the H2 line-forming region. In the power-law cloud model of de Regt et al. (2025), the cloud opacity is dominated by contributions around $\log_{10}P\sim 0.5$–1, and this estimate is robust to assumptions about patchiness. In our analysis as well, the inferred cloud top pressures consistently fall within $\log_{10}P\sim 0.5$–1, broadly agreeing with these results. On the other hand, the quantitative estimates depend on assumptions about the TTPP profile and line lists, which in turn propagate into the inferred C/O ratio as discussed in the next subsection.
VI.2 $\mathrm{C/O}$ Ratios
@@
Figure 14 compares $\mathrm{C/O}$ inferred with each CO line list configuration and with the GP TTPP model (using $\mathrm{HITEMP_{H_{2}}\,\,CO}$) to literature values for other brown dwarfs and super Jupiters. For both A and B, the choice of CO line list introduces a systematic offset of $\sim$7% in $\mathrm{C/O}$. By contrast, the assumed TTPP profile (power-law versus GP) does not yield a significant difference in $\mathrm{C/O}$. An epoch-by-epoch analysis of the Luhman 16B time series likewise reveals no statistically significant impact from cloud-cover variability on the retrieved $\mathrm{C/O}$ or molecular VMRs (Section IV.2.2). We therefore conclude that, in our framework, the principal systematic on $\mathrm{C/O}$ arises from the CO line list, whereas the TTPP shape and photometric variability are not dominant. In all cases, the C/O values for components A and B agree within $\approx 2\sigma$, consistent with the expectation that isolated brown dwarfs (and their binaries) form in a star-like manner.
@@
At face value, Figure 14 hints that brown dwarfs may exhibit higher $\mathrm{C/O}$ than super Jupiters. Although this is a collection of inhomogeneous inferences from the literature, our measurements for Luhman 16AB, together with the associated systematic-error budget, reinforce this possibility. While differences in formation pathways (i.e., gravitational instability vs. core accretion) remain a possible explanation, the contrast could also reflect differences in the efficiency of oxygen sequestration by silicate (and other) clouds, as emphasized in previous low-resolution studies (e.g., Line et al., 2015, 2017): more efficient rainout in field brown dwarfs would raise the gas-phase C/O, whereas weaker sequestration in super Jupiters would leave C/O closer to the stellar value. A decisive test will require a larger, uniformly analyzed sample of both planets and brown dwarfs, so that inter-study systematics are minimized and C/O can be compared on equal footing. Building on such a uniform baseline, joint analyses of high-resolution spectra (to constrain TTPP) and JWST coverage (to probe silicate or other cloud composition and vertical structure) can further mitigate cloud–TTPP–chemistry degeneracies, thereby providing a more direct test of the efficiency of oxygen sequestration by rainout. We return to the prospects and implementation of such joint retrievals in Section VI.4.
@@
VI.3 Usefulness and Caveats of the GP TTPP Model
For both simulations and real data, the GP TTPP model captures structures that are more complex than a simple power-law and yields a more faithful assessment of uncertainties. In the present application, these differences have only a modest impact on bulk parameters such as $\mathrm{C/O}$. However, as discussed in Section V.3, comparisons between retrieved molecular abundances and chemical-equilibrium predictions are sensitive to how accurately TTPP uncertainties are estimated; in this context, the GP framework is important because it enables predictive uncertainties to faithfully reflect the spectra’s information content, remaining conservative where the data are weakly informative. The GP approach is also sensitive to residual model–data mismatches and thus shows promise for diagnosing deficiencies in current models, potentially pointing to additional physics.
On the other hand, as indicated in Section V.3, degeneracies between the TTPP shape and cloud opacity can also allow nonphysical temperature structures if left unconstrained. Mitigations include imposing weak physical regularization (e.g., monotonicity or stability constraints) on flexible TTPP profiles and/or extending the analysis to broader wavelength coverage where cloud signatures are stronger. The latter strategy can help balance flexibility with physical plausibility; see also Section VI.4.
Figure 15: Comparison between the medium-resolution spectra predicted by extrapolating our high-resolution retrievals and the JWST observations. Shaded bands show the 90% credible intervals of the predicted spectra computed with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a power-law TTPP profile (blue: Luhman 16A; orange: Luhman 16B). Points show the corresponding JWST medium-resolution spectra from Biller et al. (2024); Chen et al. (2025) (blue: A; orange: B).
VI.4 Cloud Degeneracies and Joint Analysis with JWST
@@
We gratefully acknowledge Ian J. M. Crossfield for providing the reduced VLT/CRIRES spectra of Luhman 16AB. Based on observations collected at the European Southern Observatory under ESO programme(s) 291.C-5006(A). We also thank the anonymous reviewer for providing comments that helped improve the manuscript. This study was supported by JSPS KAKENHI Grant Numbers 24KJ1555 (H.Y.), 21H04998 (K.M., Y.K., and H.K.), 22H05150 (Y.K.), 25K17429 (Y.K.), 23H01224 (Y.K. and H.K.), and 23H00133 (H.K.). We used ChatGPT (OpenAI, GPT-5, Sept 2025) for language editing. No part of the scientific analysis was performed by ChatGPT. The authors are responsible for the content and interpretation.
Table 4 provides the literature values, together with the corresponding references, for the objects shown in Figure 1.
@@
Gl 229 B FBD Y $0.77\pm 0.01$ $71.47^{+0.57}_{-0.61}$ccGl 229 B has been reported to be a binary, based on high–angular-resolution observations with GRAVITY/VLTI together with radial-velocity monitoring with CRIRES+/VLT (Xuan et al., 2024b). $850\pm 50$ (6), (7)
@@
Note. — References—(1) de Regt et al. (2024); (2) Xuan et al. (2022); (3) Xuan et al. (2024c); (4) Hsu et al. (2024); (5) Currie et al. (2020); (6) Kawashima et al. (2025); (7) Calamari et al. (2022); (8) Wang et al. (2022); (9) Mulder et al. (2025); (10) Schlawin et al. (2017); (11) Zapatero Osorio et al. (2014); (12) Filippazzo et al. (2015); (13) Costes et al. (2024); (14) Franson et al. (2022); (15) Johnson-Groh et al. (2017); (16) González Picos et al. (2024); (17) Scholz et al. (2005); (18) Cooper et al. (2024); (19) Morales-Gutiérrez et al. (2021); (20) Zhang et al. (2021); (21) Vos et al. (2022); (22) Zhang et al. (2024b); (23) Currie et al. (2023); (24) Landman et al. (2024); (25) Lacour et al. (2021); (26) Chilcote et al. (2017); (27) González Picos et al. (2025); (28) Schwarz et al. (2016); (29) Zhang et al. (2024a); (30) Wood et al. (2023); (31) Xuan et al. (2024a); (32) Gandhi et al. (2025); (33) Bonnefoy et al. (2010)